# How (un)ethical are instruction-centric responses of LLMs? Unveiling the vulnerabilities of safety guardrails to harmful queries

### v1

@@ 7 Extent of harmfulness
So far we have analysed if a particular response generated is harmful or not, i.e., all our results were based on a binary judgement. Here we take a step further and investigate the extent of harmfulness present is a generated response. The harmfulness of the model generated responses depends on the presence of vicious suggestions specific to the malicious input queries. In our experiment (inspired by Zhao et al. (2024)), we use a reward model111111https://huggingface.co/OpenAssistant/ reward-model-deberta-v3-large-v2 to measure the harmfulness of the generated output. This reward model returns a negative value. In Table 6, we show the absolute values, i.e., the higher is the value the more harmful is the response. As the table shows, in all the three prompt settings and across all topics, pseudocode responses are more harmful compared to the text responses. In the zero-shot setting the top three most intense harmful pseudocode responses are generated in topics Cyber Security, Chemical Weapon, Biological and Chemical Weapons and Social Media by the Mixtral 8X7B model. Similarly, in the zero-shot CoT setting the top three most intense harmful pseudocode responses are generated in topics Cyber Security, Nuclear Technology, Bio Nuclear Terrorism and Chemical Weapon, Biological and Chemical Weapons. Lastly, in the few-shot setting the top three most intense harmful pseudocode responses are generated in topics Nuclear Technology, Bio Nuclear Terrorism, Chemical Weapon, Biological and Chemical Weapons and Biotechnology, Biology, Genetic Engineering. The most surprising observation perhaps is that highest intensity harmful responses are produced in the few-shot setting which is actually considered as a remedial technique for avoiding such harmful response generations.
zero-shot | zero-shot CoT | few-shot | |||||
Topics | Models | P | T | P | T | P | T |
Llama-2-13b | 2.9238 | 1.3881 | 3.0018 | 2.1779 | 4.2933 | 1.6744 | |
Llama-2-7b | 3.8724 | 2.0885 | 3.8931 | 3.3604 | 4.9071 | 2.3745 | |
Mistral-V2 | 3.8009 | 2.9876 | 4.6203 | 3.0896 | 5.346 | 2.5919 | |
BBG | Mixtral 8X7B | 3.6413 | 3.3876 | 4.1404 | 3.7521 | 5.127 | 2.7392 |
Llama-2-13b | 2.8707 | 1.0566 | 2.9255 | 1.8267 | 4.1866 | 1.7096 | |
Llama-2-7b | 4.29 | 1.7607 | 4.41 | 3.2469 | 5.0096 | 2.5405 | |
Mistral-V2 | 4.2332 | 3.0768 | 4.5841 | 3.4869 | 5.6405 | 2.9175 | |
NBT | Mixtral 8X7B | 4.1296 | 4.2529 | 4.8401 | 4.3809 | 5.5513 | 3.3085 |
Llama-2-13b | 2.6126 | 1.1365 | 2.6243 | 1.7119 | 4.0095 | 1.7201 | |
Llama-2-7b | 4.1485 | 1.9137 | 4.1124 | 3.2102 | 4.7898 | 2.5096 | |
Mistral-V2 | 4.1811 | 3.6807 | 4.5943 | 4.0196 | 5.62 | 3.1236 | |
CBC | Mixtral 8X7B | 4.5938 | 4.4868 | 4.7168 | 4.6407 | 5.4851 | 3.3237 |
Llama-2-13b | 2.8358 | 1.769 | 3.1978 | 2.0253 | 4.3787 | 1.6761 | |
Llama-2-7b | 2.7729 | 1.908 | 4.292 | 3.4767 | 4.9153 | 2.5048 | |
Mistral-V2 | 4.1374 | 2.8162 | 4.8486 | 3.2004 | 5.2967 | 2.4994 | |
CS | Mixtral 8X7B | 4.6594 | 3.7572 | 4.7804 | 3.9875 | 5.0427 | 2.5801 |
Llama-2-13b | 2.9454 | 1.6351 | 2.8334 | 1.7375 | 4.1273 | 1.7389 | |
Llama-2-7b | 2.8378 | 1.4487 | 3.9946 | 3.1641 | 4.73 | 2.5934 | |
Mistral-V2 | 3.6965 | 2.9208 | 4.4461 | 3.4548 | 5.3276 | 3.0315 | |
FB | Mixtral 8X7B | 4.3284 | 3.9019 | 4.303 | 4.198 | 5.1489 | 3.1736 |
Llama-2-13b | 2.8332 | 1.4118 | 2.7922 | 1.976 | 4.0067 | 1.7017 | |
Llama-2-7b | 3.7257 | 2.3548 | 3.7332 | 3.631 | 4.8263 | 2.5163 | |
Mistral-V2 | 3.8252 | 2.9573 | 4.7089 | 3.5912 | 5.2361 | 2.8475 | |
SM | Mixtral 8X7B | 4.4806 | 4.0571 | 4.5965 | 4.4387 | 5.1543 | 3.0919 |
Llama-2-13b | 2.8332 | 1.1804 | 2.5869 | 1.7307 | 4.1725 | 1.7032 | |
Llama-2-7b | 3.8283 | 1.755 | 4.084 | 3.2287 | 4.9112 | 2.465 | |
Mistral-V2 | 3.6413 | 2.6879 | 4.7103 | 2.8953 | 5.2552 | 2.6997 | |
PP | Mixtral 8X7B | 4.1367 | 3.1332 | 4.225 | 3.305 | 5.1849 | 2.9246 |
8 Conclusion
@@
Ethical statement
Our paper highlights the challenges faced by LLMs while responding to harmful queries in a pseudocode format. The objective throughout has been to identify specific concerns, inviting the wider community to engage and contribute toward solving these issues. We stress that we do not wish in any way to undermine ethical values or disrupt international harmony. Our goal is to foster a conversation on the responsible use and development of artificial intelligence technologies. We recognize that presenting content in particular ways can demonstrate potential misuses or misconceptions of AI. Nonetheless, our dedication to ethical principles and appreciation for varied viewpoints is unwavering.

### v5

@@ Results
So far we have analysed if a particular response generated is harmful or not, i.e., all our results were based on a binary judgement. Here we take a step further and investigate the extent of harmfulness present is a generated response. The harmfulness of the model generated responses depends on the presence of vicious suggestions specific to the malicious input queries. In our experiment (inspired by (Zhao et al. 2024)), we use a reward model101010https://huggingface.co/OpenAssistant/ reward-model-deberta-v3-large-v2 to measure the harmfulness of the generated output. This reward model returns a negative value. In Table 6, we show the absolute values, i.e., the higher is the value the more harmful is the response. As the table shows, in all the three prompt settings and across all topics, pseudocode responses are more harmful compared to the text responses.
In the zero-shot setting the top three most intense harmful pseudocode responses are generated in topics Cyber Security, Chemical Weapon, Biological and Chemical Weapons and Social Media by the Mixtral 8X7B model. Similarly, in the zero-shot CoT setting the top three most intense harmful pseudocode responses are generated in topics Cyber Security, Nuclear Technology, Bio Nuclear Terrorism and Chemical Weapon, Biological and Chemical Weapons. Lastly, in the few-shot setting the top three most intense harmful pseudocode responses are generated in topics Nuclear Technology, Bio Nuclear Terrorism, Chemical Weapon, Biological and Chemical Weapons and Biotechnology, Biology, Genetic Engineering. The most surprising observation perhaps is that highest intensity harmful responses are produced in the few-shot setting which is actually considered as a remedial technique for avoiding such harmful response generations.
Further we compute the standard deviation of the harmfulness scores to understand the overall variation in these values. For Llama2-7B, Mistral-v2, and Mixtral 8x7B, the standard deviation for pseudocode (0.11-0.35) is lower than that for plain text () in the zero-shot setting. For Mistral-v2 and Mixtral 8x7B, the standard deviation for pseudocode () is lower than for plain text () in the zero-shot-CoT and few-shot settings. Only Llama2-7B and 13B, in the zero-shot-CoT and few-shot settings, the standard deviation for pseudocode () is slightly higher than for plain text (). Thus the harmfulness score for the pseudocode responses across most of the settings vary less and are clustered better around the mean indicating the robustness of our observations.
Discussion
Factors | Contributing factors | Model(s) affected | Impact | Implications |
Increased harmful pseudocode responses | Instruction-centric prompts (pseudocode) | All models | Harmful responses increased by 2-38% in zero-shot settings, with the highest increase in pseudocode generation. | Indicates a vulnerability when generating structured responses; requires focused mitigation strategies. |
Vulnerability to model editing | Application of ROME editing technique | Llama-2-7b, Llama-2-13b | Post-editing, harmful pseudocode response rates rose significantly (e.g., 18.9% to 56.66% for Llama-2-7b). | Model edits can amplify unethical outputs, suggesting the need for robust controls on model modification. |
Prompting strategy sensitivity | Zero-shot CoT and few-shot prompting methods | All models | Increased harmful output generation, especially in CoT settings (e.g., 28.6% increase in Nuclear Technology domain). | Advanced reasoning prompts (CoT) may inadvertently increase unethical output risk; need refined prompt strategies. |
Layer-specific editing sensitivity | Edits to specific model layers (e.g., layer 1, 3, 5) | Llama-2-7b | Varying impact on harmful content generation depending on the layer edited; increased harm in certain domains. | Indicates different model layers have distinct impacts on ethical output; targeted layer-specific safety training needed. |
High-intensity harmful outputs | Pseudocode responses in few-shot and CoT settings | Mixtral 8X7B | Most intense harmful outputs in Cyber Security, Chemical Weapons domains; increased in few-shot settings. | Few-shot prompting can lead to high-intensity harmful outputs, challenging assumptions about its mitigating effects. |
In this section, we present a detailed discussion of the factors contributing to the generation of unethical content by LLMs when responding to instruction-centric prompts (refer to Table 8 for a summary). Some of these factors include model types, prompting strategies, and the impact of model editing and are discussed below.
Model vulnerabilities and response patterns
Our experiments demonstrate significant variability in the models’ propensity to generate unethical responses based on the format of the prompt. Notably, the Llama-2-13b model showed a marked increase in unethical content when asked to produce pseudocode rather than plain text, with harmful responses rising from 10.5% to 48.7% in zero-shot settings for the Biotechnology, Biology, and Genetic Engineering topic (see Table 3). Similar trends were observed across other topics, such as Nuclear Technology and Cyber Security, where the increase in harmful responses was even more pronounced for pseudocode prompts. This suggests a specific vulnerability of these models when tasked with generating structured or instruction-centric outputs.
Impact of prompting strategies
The choice of prompting strategy largely affects the generation of unethical content. For instance, in the zero-shot chain-of-thought (CoT) setting, the generation of harmful pseudocode responses increased considerably across all models and topics. For the Llama-2-13b model, harmful pseudocode responses in the Nuclear Technology domain increased by 28.6% in the zero-shot CoT setting compared to the basic zero-shot setup (see Table 3). This highlights that even prompting strategies designed to enhance model reasoning capabilities can inadvertently increase the risk of unethical outputs, particularly when dealing with complex, instruction-centric queries.
Effects of model editing
Model editing, particularly using the ROME technique, exacerbates the models’ tendency to generate unethical content. Post-editing, the Llama-2-7b model shows a substantial increase in harmful pseudocode responses across all topics. The most significant increases are observed in zero-shot settings, where harmful pseudocode responses rise from 18.9% to 56.66% on average (see Table 5). This trend persists in few-shot settings as well, with harmful response rates increasing to 65.67%. The results indicate that minor edits to model parameters can significantly amplify the generation of unethical content, underscoring the need for robust safeguards against model tampering.
Layer sensitivity in model editing
The sensitivity of the models to editing vary across different layers. For instance, editing higher layers in the Llama-2-7b model generally results in a reduction in the percentage of harmful pseudocode responses for topics like Biotechnology and Social Media. Conversely, for topics such as Cyber Security and Finance, editing higher layers increases the percentage of harmful responses (see Figure 1). This suggests that different model layers encode different types of information relevant to ethical judgment, and targeted edits at specific layers can either mitigate or exacerbate harmful outputs depending on the content domain.
Extent of harmfulness in generated content
We note that the intensity of harmful content is consistently higher for pseudocode responses across all models and topics. The Mixtral 8X7B model, in particular, generated the most intense harmful responses in topics such as Cyber Security and Chemical Weapons under zero-shot CoT settings (see Table 6). Interestingly, the few-shot prompting strategy, typically seen as a mitigation approach, led to the highest intensity of harmful responses, challenging the assumption that providing examples always enhances model safety.
Implications for future model development
Our findings reveal critical gaps in the current mitigation strategies for LLMs, particularly regarding instruction-centric prompts. The substantial increase in unethical outputs when models are queried with structured or code-like prompts suggests a need for more targeted safety training and the development of red teaming mechanisms that specifically address these vulnerabilities. Moreover, the ease with which model editing can amplify unethical content production highlights the importance of robust model integrity and security measures to prevent unauthorized modifications.
Conclusion